The MPS for the product shown calls for 75 units to be produced in weeks 2 and 4. Item B has a lead time of 1 week, the batch size is 200, and there are currently 100 units of inventory. A planned order release for B will be needed in week(s) ________.

  • a. 1
  • b. 2
  • c. 3
  • d. 1 and 2
  • e. 1 and 3A B (3) С D (2)
Answer #1

Answer: e. 1 and 3

Explanation:

MPS for Product A in Week 2 = 75

Gross Requirement of B in Week 2 = 75*3 = 225

Net Requirement of B in Week 2 = Gross Requirement - Current Inventory = 225 - 100 = 125

Planned Order Release of B in week 1 = 200 (As Lead time = 1 week and Lot Size = 200)

Projected On Hand for Week 2 = Planned Order Receipt - Net Requirement = 200 - 125 =75

MPS for Product A in Week 4 = 75

Gross Requirement of B in Week 4 = 75*3 = 225

Net Requirement of B in Week 4 = Gross Requirement - Current Inventory = 225 - 75 = 150

Planned Order Release of B in week 3 = 200 (As Lead time = 1 week and Lot Size = 200)

Projected On Hand for Week 4 = Planned Order Receipt - Net Requirement = 200 - 150 =50